Twice Martina's age decreased by Mario's age is 6. Three times Martina's age increased by twice Mario's is 23. What are the two numbers?
Answer by stanbon(75887)

Twice Martina's age decreased by Mario's age is 6. Three times Martina's age increased by twice Mario's is 23. What are the two numbers?
--------------------
r - t = 6
3r + 2t = 23
-------------------
Modify for elimination:
2r - 2t = 12
3r + 2t = 23
-------------------
Add and solve for "r"::
5r = 35
r = 7 (Martina's age)
-----
Solve for "t"::
r - t = 6
7 - t = 6
t = 1 (Mario's age)
-------------
Cheers,
Stan H.
--------------